# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v3

@@ III-B Real-Scene Based Simulators
Table III summarizes all the simulators based on the real scenarios mentioned above. Sapien is quite special among these, it is completely designed for simulating interactions with joint objects (such as doors, cabinets, and drawers); Virtualhome’s unique environment graph makes it widely used for high-level embodied planning based on the natural-language-described environment; AI2Thor provides a wealth of interactive scenes, but these interactions, like Virtualhome, are only implemented through scripts and do not have real physical interactions. However, such a design is sufficient for embodied tasks that do not focus on fine-grained interactions; both iGibson and TDW provide fine-grained embodied control and highly simulated physical interactions, among which iGibson provides more abundant and realistic large-scale scenes, which makes iGibson suitable for more complex and long-term mobile operation tasks, while TDW provides users with higher freedom, users can freely expand the scene, and TDW’s unique audio and flexible, fluid simulation make it irreplaceable in related simulation scenarios; Matterport3D, as a basic 2D-3D visual dataset, is widely used and extended in the current embodied AI benchmark; in Habitat, the embodied agent lacks interaction capabilities, but Habitat’s huge indoor scenes and easy-to-use interfaces, open framework, make it highly regarded in the field of embodied navigation.
@@ IV Embodied Perception
The “north stars” of the future of visual perception is embodied-centric visual reasoning and social intelligence [76]. Unlike merely recognizing objects in images [77, 78, 79], agent with embodied perception must move in the physical world and interact with the environment. This requires a deeper understanding of 3D space and dynamic environments. Embodied perception requires visual perception and reasoning, understanding the 3D relations within a scene, and predicting and performing complex tasks based on visual information.
@@ IV-B2 One-stage 3D Visual Grounding methods
As shown in Fig. 8 (c), in contrast to two-stage 3D VG methods, one-stage 3D VG methods integrate object detection and feature extraction guided by language queries, making it easier to locate objects relevant to the language.
@@ IV-C Visual Language Navigation
SR (Success Rate), TL (Trajectory Length), and SPL (Success Weighted by Path Length) are the most commonly used metrics in VLN. Among them, SR directly reflects the navigation performance of the embodied agent, TL reflects the navigation efficiency, and SPL combines both to indicate the overall performance of the embodied agent.
Below, we will introduce VLN divided into two parts: datasets and methods.
@@ IV-C1 Datasets
Room to Room [65] is a VLN dataset created by Anderson et al. based on Matterport3D. In R2R, embodied agents navigate according to step-by-step instructions, choosing the next adjacent navigation graph node to advance based on visual observations until they reach the target location. Embodied agents need to dynamically track progress to align the navigation process with fine-grained instructions. Room-for-Room [162] extends the paths in R2R to longer trajectories, requires higher for the long-distance instruction and history alignment capabilities of embodied agents. VLN-CE [163] extends R2R and R4R to continuous environments, embodied agents can move freely in the scene. This makes the action decision of embodied agents more difficult. Different from the above datasets based on indoor scenes, Chen et al. created the TOUCHDOWN dataset [164] based on Google Street View. In TOUCHDOWN, embodied agents follow instructions to navigate in the street view rendering simulation of New York City to find the specified object.
@@
Shridhar et al. [168] proposed the ALFRED dataset based on the AI2-THOR simulator. In ALFRED, embodied agents need to understand environmental observations and complete household tasks in an interactive environment according to coarse-grained and fine-grained instructions. The task of OVMM [169] is to pick any object in any unseen environment and place it in a specified location. Embodied agents need to locate the target object in the home environment, navigate and grab it, and then navigate to the target location to put down the object. OVMM provides a simulation based on Habitat and a framework for implementation in the real world. Behavior-1K [170] is a dataset based on human needs, comprising 1,000 long-sequence, complex, skill-dependent daily tasks, which are designed within OmniGibson, an extension of the iGibson simulation environment. Embodied agents need to complete long-span navigation-interaction tasks which may contain thousands of low-level action steps based on visual information and natural language instructions. Such complex tasks pose higher requirements for the understanding and memory of embodied agents.
@@ IV-C2 Method
VLN has made great strides in recent years. With the astonishing performance of large models in various fields, the direction and focus of VLN research have been profoundly influenced. But overall, the research methods of VLN can still be roughly divided into two directions: Memory-Understanding Based, and Future-Prediction Based.
@@ V-B Embodied Grasping
Embodied interaction, in addition to engaging in question-answering interactions with humans, can also involve performing operations based on human instructions, such as grasping and placing objects, thereby completing interactions among the robots, humans and objects. Embodied grasping requires comprehensive semantic understanding, scene perception, decision-making, and robust control planning. The embodied grasping methods integrate traditional robotic kinematic grasping with large models such as large language models (LLMs) [286] and vision-language foundation models [16], which enables agents to perform grasping tasks under multi-sensory perceptions, including visual active perception, language understanding and reasoning. Figure 12 (b) illustrates an overview of human-agent-object interaction, where the agent accomplishes embodied grasping tasks.
@@ V-B3 Language-guided grasping
In both cases, the integration of semantic understanding with spatial and logical reasoning enables the agent to perform complex grasping tasks effectively and accurately. Figure 12 (a) depicts instances of various types of language-guided grasping tasks described above.
@@ VI Embodied Agent
An agent is defined as an autonomous entity capable of perceiving its environment and acting to achieve specific objectives. Initially, Symbolic Agents, rooted in symbolic reasoning, and Reactive Agents, known for their rapid responsiveness, were widely utilized. However, these agents were limited in handling complex strategies. Learning-based agents were subsequently developed to mitigate this limitation, yet they remained inadequate for large-scale real-world problems. Recent advancements in MLMs have further expanded the application of agents to practical scenarios. When these MLM-based agents are embodied in physical entities, they can effectively transfer their capabilities from virtual space to the physical world, thereby becoming embodied agents [304].
@@ VI-B3 Planning utilizing the VLMs
Compared to converting environmental information into text using external visual models, VLM models can capture visual details in latent space, particularly contextual information that is difficult to represent with object labels, as per Lecun’s insights on world models. VLMs can discern rules underlying visual phenomena; for instance, even if a towel is not visible in the environment, it can be inferred that the towel might be stored in a cabinet. This process essentially demonstrates how abstract visual features and structured textual features can be more effectively aligned in latent space.
In the EmbodiedGPT framework, the Embodied-Former module aligns embodied, visual, and textual information, effectively considering the agent’s state and environmental information during task planning [310]. Unlike EmbodiedGPT, which directly uses third-person perspective images, LEO [347] encodes 2D egocentric images and 3D scenes into visual tokens. This method effectively perceives 3D world information and executes tasks accordingly. Similarly, the EIF-Unknow model utilizes Semantic Feature Maps extracted from Voxel Features as visual tokens, which are input along with text tokens into a trained LLaVA model for task planning [348]. Furthermore, embodied multimodal foundation models, or VLA models, have been extensively trained with large datasets in studies like the RT series [13, 308] and PaLM-E [307] to achieve alignment of visual and textual features in embodied scenarios. Additionally, innovative research such as Matcha [349] explores the integration of large language models (LLMs) with multimodal perception modules (e.g., weight, touch, sound) to enhance robot planning capabilities. For example, when executing the task “pick up a plastic block,” the robot can assess its weight by lifting it, determine if the sound it makes when tapped matches expected plastic sounds, or use tactile feedback to sense the block’s hardness or softness. Moreover, the VLP [350] model enhances planning by considering historical information and predicting future visual states. Once an LLM generates an action, the video model can simulate multiple potential video representations for each action. These video models then serve as heuristic functions to evaluate the proximity of each action to the goal, thereby improving planning effectiveness.
@@ VI-C Embodied Action Planning
Section VI-B discusses the definitions and differences between task planning and action planning. It is evident that action planning must address real-world uncertainties because the granularity of subtasks provided by task planning is insufficient to guide agents in efficient environmental interaction. Generally, agents can approach action planning in two ways: first, by using pre-trained embodied perception models and embodied intervention models as tools, incrementally completing the subtasks specified by task planning through APIs; second, by utilizing the VLA model’s inherent capabilities to derive action planning. Furthermore, the results of the action planner’s execution are fed back to the task planner to adjust and improve task planning.
@@ VII-A3 Knowledge-driven Methods
Knowledge-driven world models inject artificially constructed knowledge into the models, endowing them with world knowledge. This method has shown broad application potential in the field of embodied AI. For example, in the real2sim-to-real approach[370], real-world knowledge is used to build physics-compliant simulators, which are then used to train robots, enhancing model robustness and generalization capabilities. Additionally, artificially constructing common sense or physics-compliant knowledge and applying them to generative models or simulators is a common strategy (e.g., ElastoGen[364], Liu et al.[365], Lionel et al. [363]). This approach imposes more physically accurate constraints on the model, enhancing its reliability and interpretability in generation tasks. These constraints ensure the model’s knowledge is both accurate and consistent, reducing uncertainty during training and application. Some approaches combine artificially created physical rules with large language models (LLMs) or large multimodal models (LMMs). By leveraging the commonsense capabilities of LLMs and LMMs, these approaches (e.g., Holodeck[73], LEGENT[366]) generate diverse and semantically rich scenes through automatic spatial layout optimization. This greatly advances the development of general-purpose embodied agents by training them in novel and diverse environments.
@@ VII-B1 Real-World Data
Training large, high-capacity models on high-volume, rich datasets has demonstrated remarkable capabilities and significant success in effectively addressing downstream applications. For instance, large language models such as ChatGPT, GPT-4, and LLaMA have not only excelled in the field of NLP but have also provided excellent problem-solving capabilities for downstream tasks. Therefore, is it possible to train an embodied large model in the robotics field, one that possesses strong generalization capabilities through training and can adapt to new scenarios and robotic tasks? This requires a large volume of embodied datasets to provide data for model training. Open X-Embodiment [309] is an embodied dataset from 22 different robots, showcasing 527 skills and 160,266 tasks. The data they collected consisted of authentic demonstration data from robots, obtained by recording the process of executing operations. This primarily focused on domestic and kitchen settings, involving items such as furniture, food, and tableware. The operations mainly centered around pick-and-place tasks, with a small portion involving more complex maneuvers. The high-capacity model RT-X trained on this dataset demonstrated excellent transfer capabilities. UMI [371] proposed a data collection and policy learning framework. They designed a handheld gripper and an elegant interface for data collection, enabling portable, low-cost, and information-rich data collection for challenging bimanual and dynamic demonstration data. By simply modifying the training data, the robot can achieve zero-shot generalizable, bimanual, and precise tasks. Mobile ALOHA [372] is a low-cost full-body mobile manipulation system. It can be used to collect task data for bimanual operations under full-body mobility, such as frying shrimp and serving dishes. Training agents with data collected by this system and static ALOHA can improve the performance of mobile manipulation tasks. Such agents can serve as home assistants or work assistants. In human-agent collaboration [373], humans and agents learn together during data collection, reducing human workload, accelerating data acquisition, and improving data quality. Specifically, in an embodied scenario, during data collection, humans provide initial action inputs. Subsequently, the agent refines these actions through iterative perturbation and denoising processes, gradually optimizing them to produce precise, high-quality operational demonstrations. This entire process can be summarized as follows: humans contribute intuition and diversity in operations, while agents handle optimization and stability, reducing reliance on operators, enabling execution of more complex tasks, and gathering higher-quality data.
@@ VII-B3 Sim-to-Real Paradigms
Real2Sim2real [376] enhances imitation learning in real-world scenarios by leveraging reinforcement learning (RL) trained in a “digital twin” simulation environment. This method involves strengthening strategies through extensive RL within the simulation, followed by transferring these strategies to the real world to address data scarcity and achieve effective robotic operational imitation learning. Initially, NeRF and VR are used for scene scanning and reconstruction, and the constructed scene assets are imported into the simulator to achieve real-to-simulation fidelity. Subsequently, RL in the simulation fine-tunes the initial strategies derived from sparse expert demonstrations collected in the real world. Finally, the refined strategies are transferred to real-world settings.
@@
TRANSIC [377] narrows the sim-to-real gap by enabling real-time human intervention to correct robot behaviors in real-world scenes. It enhances sim-to-real transfer performance through several steps: first, robots are trained using RL to establish foundational strategies within a simulation environment. Then, these strategies are implemented on real robots, with humans intervening and correcting behaviors in real-time via remote control when errors occur. The data collected from these interventions are used to train a residual policy. Integrating both foundational and residual policies ensures smoother trajectories in real-world applications following sim-to-real transfer. This approach significantly reduces the need for real-world data collection, thereby mitigating the burden while achieving effective sim-to-real transfer.
@@
Domain Randomization [378, 379, 380] enhances the generalization of models trained in simulated environments to real-world scenarios by introducing parameter randomization during simulation. While both simulated and real environments involve perception through camera-acquired visual images, differences such as object friction and gloss make accurate simulation challenging. Randomizing parameters during simulation training covers a wide range of conditions, potentially encompassing variations that might occur in real-world settings. This approach boosts the robustness of trained models, enabling deployment from simulation to real environments.
@@
System Identification [381, 382] constructs an accurate mathematical model of physical scenes in real-world environments, encompassing parameters such as dynamics and visual rendering. It aims to make simulation environments closely resemble real-world settings, facilitating smooth transitions of models trained in simulation to real environments.
@@
Lang4sim-to-real [383] uses natural language as a bridge to address the sim-to-real gap by using textual descriptions of images as a cross-domain unified signal. This approach aids in learning domain-invariant image representations, thereby improving generalization performance across simulation and real environments. Initially, an encoder is pretrained on image data annotated with cross-domain language descriptions. Subsequently, using the learned domain-invariant representations, a multi-domain, multi-task language-conditioned behavioral cloning policy is trained. This method compensates for the scarcity of real-world data by leveraging additional information from abundant and inexpensive simulated data, thereby enhancing sim-to-real transfer.

### v4

@@ III-B Real-Scene Based Simulators
Table III summarizes all the simulators based on the real scenarios mentioned above. Sapien stands out for its design, specifically tailored to simulate interactions with joint objects like doors, cabinets, and drawers. VirtualHome is notable for its unique environment graph, which facilitates high-level embodied planning based on natural language descriptions of environments. While AI2Thor offers a wealth of interactive scenes, these interactions, similar to those in VirtualHome, are script-based and lack real physical interactions. This design suffices for embodied tasks not requiring fine-grained interactions. Both iGibson and TDW provide fine-grained embodied control and highly simulated physical interactions. iGibson excels in offering abundant and realistic large-scale scenes, making it suitable for complex and long-term mobile operations, whereas TDW allows greater user freedom in scene expansion and features unique audio and flexible fluid simulations, making it indispensable for related simulation scenarios. Matterport3D, a foundational 2D-3D visual dataset, is widely used and extended in embodied AI benchmarks. Although the embodied agent in Habitat lacks interaction capabilities, its extensive indoor scenes, user-friendly interfaces, and open framework make it highly regarded in embodied navigation.
@@ IV Embodied Perception
The “north stars” of the future of visual perception is embodied-centric visual reasoning and social intelligence [76]. Unlike merely recognizing objects in images, agent with embodied perception must move in the physical world and interact with the environment. This requires a deeper understanding of 3D space and dynamic environments. Embodied perception requires visual perception and reasoning, understanding the 3D relations within a scene, and predicting and performing complex tasks based on visual information.
@@ IV-B2 One-stage 3D Visual Grounding methods
In Fig. 8 (c), in contrast to two-stage 3D VG methods, one-stage 3D VG methods integrate object detection and feature extraction guided by language queries, making it easier to locate objects relevant to the language.
@@ IV-C Visual Language Navigation
SR (Success Rate), TL (Trajectory Length), and SPL (Success Weighted by Path Length) are the most commonly used metrics in VLN. Among them, SR directly reflects the navigation performance of the embodied agent, TL reflects the navigation efficiency, and SPL combines both to indicate the overall performance of the embodied agent. Below, we introduce VLN in two parts: datasets and methods.
@@ IV-C1 Datasets
Room to Room (R2R) [65] is a VLN dataset based on Matterport3D. In R2R, embodied agents navigate according to step-by-step instructions, choosing the next adjacent navigation graph node to advance based on visual observations until they reach the target location. Embodied agents need to dynamically track progress to align the navigation process with fine-grained instructions. Room-for-Room [159] extends the paths in R2R to longer trajectories, which requires stronger long-distance instruction and history alignment capabilities of embodied agents. VLN-CE [160] extends R2R and R4R to continuous environments, embodied agents can move freely in the scene. This makes the action decision of embodied agents more difficult. Different from the above datasets based on indoor scenes, TOUCHDOWN dataset [161] is created based on Google Street View. In TOUCHDOWN, embodied agents follow instructions to navigate in the street view rendering simulation of New York City to find the specified object.
@@
ALFRED dataset [165] is based on the AI2-THOR simulator. In ALFRED, embodied agents need to understand environmental observations and complete household tasks in an interactive environment according to coarse-grained and fine-grained instructions. The task of OVMM [166] is to pick any object in any unseen environment and place it in a specified location. Agents need to locate the target object in the home environment, navigate and grab it, and then navigate to the target location to put down the object. OVMM provides a simulation based on Habitat and a framework for implementation in the real world. Behavior-1K dataset [167] is based on human needs, comprising 1,000 long-sequence, complex, skill-dependent daily tasks, which are designed within OmniGibson, an extension of the iGibson simulation environment. Agents need to complete long-span navigation-interaction tasks which contain thousands of low-level action steps based on visual information and language instructions. These complex tasks requires strong capabilities of understanding and memory.
@@ IV-C2 Method
VLN has made great strides recently with the astonishing performance of LLMs, the direction and focus of VLN have been profoundly influenced. Nevertheless, the VLN methods can be divided into two directions: Memory-Understanding Based and Future-Prediction Based.
@@ V-B Embodied Grasping
Embodied interaction, in addition to engaging in question-answering interactions with humans, can also involve performing operations based on human instructions, such as grasping and placing objects, thereby completing interactions among the robots, humans and objects. Embodied grasping requires comprehensive semantic understanding, scene perception, decision-making, and robust control planning. The embodied grasping methods integrate traditional robotic kinematic grasping with large models such as LLMs [283] and vision-language foundation models [16], which enables agents to perform grasping tasks under multi-sensory perceptions, including visual active perception, language understanding and reasoning. Figure 12 (b) illustrates an overview of human-agent-object interaction, where the agent accomplishes embodied grasping tasks.
@@ V-B3 Language-guided grasping
In both cases, the integration of semantic understanding with spatial and logical reasoning enables the agent to perform complex grasping tasks effectively and accurately. Figure 12 (a) depicts various types of language-guided grasping tasks.
@@ VI Embodied Agent
An agent is defined as an autonomous entity capable of perceiving its environment and acting to achieve specific objectives. Recent advancements in MLMs have further expanded the application of agents to practical scenarios. When these MLM-based agents are embodied in physical entities, they can effectively transfer their capabilities from virtual space to the physical world, thereby becoming embodied agents [301].
@@ VI-B3 Planning utilizing the VLMs
Compared to converting environmental information into text using external visual models, VLM models can capture visual details in latent space, particularly contextual information that is difficult to represent with object labels. VLMs can discern rules underlying visual phenomena; for instance, even if a towel is not visible in the environment, it can be inferred that the towel might be stored in a cabinet. This process essentially demonstrates how abstract visual features and structured textual features can be more effectively aligned in latent space.
In EmbodiedGPT [307], the Embodied-Former module aligns embodied, visual, and textual information, effectively considering the agent’s state and environmental information during task planning. Unlike EmbodiedGPT, which directly uses third-person perspective images, LEO [344] encodes 2D egocentric images and 3D scenes into visual tokens. This method effectively perceives 3D world information and executes tasks accordingly. Similarly, the EIF-Unknow model utilizes Semantic Feature Maps extracted from Voxel Features as visual tokens, which are input along with text tokens into a trained LLaVA model for task planning [345]. Furthermore, embodied multimodal foundation models, or VLA models, have been extensively trained with large datasets in studies like the RT series [13, 305] and PaLM-E [304] to achieve alignment of visual and textual features in embodied scenarios. Additionally, innovative research such as Matcha [346] explores the integration of LLMs with multimodal perception modules (e.g., weight, touch, sound) to enhance robot planning capabilities. For example, when executing the task “pick up a plastic block,” the robot can assess its weight by lifting it, determine if the sound it makes when tapped matches expected plastic sounds, or use tactile feedback to sense the block’s hardness or softness. Moreover, the VLP [347] model enhances planning by considering historical information and predicting future visual states. Once the LLM generates an action, the video model can simulate multiple potential video representations for each action. These video models then serve as heuristic functions to evaluate the proximity of each action to the goal, thereby improving planning effectiveness.
@@ VI-C Embodied Action Planning
Section VI-B discusses the definitions and differences between task planning and action planning. It is evident that action planning must address real-world uncertainties because the granularity of subtasks provided by task planning is insufficient to guide agents in environmental interaction. Generally, agents can achieve action planning in two ways: 1) using pre-trained embodied perception and embodied intervention models as tools, incrementally completing the subtasks specified by task planning through APIs, 2) utilizing the VLA model’s inherent capabilities to derive action planning. Furthermore, the results of the action planner’s execution are fed back to the task planner to adjust and improve task planning.
@@ VII-A3 Knowledge-driven Methods
Knowledge-driven world models inject artificially constructed knowledge into the models, endowing them with world knowledge. This method has shown broad application potential in the field of embodied AI. For example, in the real2sim2real approach[367], real-world knowledge is used to build physics-compliant simulators, which are then used to train robots, enhancing model robustness and generalization capabilities. Additionally, artificially constructing common sense or physics-compliant knowledge and applying them to generative models or simulators is a common strategy (e.g., ElastoGen [361], One-2-3-45 [362], PLoT [360]). This approach imposes more physically accurate constraints on the model, enhancing its reliability and interpretability in generation tasks. These constraints ensure the model’s knowledge is both accurate and consistent, reducing uncertainty during training and application. Some approaches combine artificially created physical rules with LLMs or MLMs. By leveraging the commonsense capabilities of LLMs and LMMs, these approaches (e.g., Holodeck[73], LEGENT[363]) generate diverse and semantically rich scenes through automatic spatial layout optimization. This greatly advances the development of general-purpose embodied agents by training them in novel and diverse environments.
@@ VII-B1 Real-World Data
Training large, high-capacity models on high-volume, rich datasets has demonstrated remarkable capabilities and significant success in effectively addressing downstream applications. For instance, LLMs such as ChatGPT, GPT-4, and LLaMA have not only excelled in the field of NLP but have also provided excellent problem-solving capabilities for downstream tasks. Therefore, is it possible to train an embodied large model in the robotics field, which possesses strong generalization capabilities through training and can adapt to new scenarios and robotic tasks. This requires a large volume of embodied datasets to provide data for model training. Open X-Embodiment [306] is an embodied dataset from 22 different robots, with 527 skills and 160,266 tasks. The collected data consists of authentic demonstration data from robots, obtained by recording the process of executing operations. This dataset primarily focuses on domestic and kitchen settings, involving items such as furniture, food, and tableware. The operations are mainly centered around pick-and-place tasks, with a small portion involving more complex maneuvers. The high-capacity model RT-X trained on this dataset demonstrates excellent transfer capabilities. UMI [368] proposed a data collection and policy learning framework. They designed a handheld gripper and an elegant interface for data collection, enabling portable, low-cost, and information-rich data collection for challenging bimanual and dynamic demonstration data. By simply modifying the training data, the robot can achieve zero-shot generalizable, bimanual, and precise tasks. Mobile ALOHA [369] is a low-cost full-body mobile manipulation system. It can be used to collect task data for bimanual operations under full-body mobility, such as frying shrimp and serving dishes. Training agents with data collected by this system and static ALOHA can improve the performance of mobile manipulation tasks. Such agents can serve as home assistants or work assistants. In human-agent collaboration [370], humans and agents learn together during data collection, reducing human workload, accelerating data acquisition, and improving data quality. Specifically, in an embodied scenario, during data collection, humans provide initial action inputs. Subsequently, the agent refines these actions through iterative perturbation and denoising processes, gradually optimizing them to produce precise, high-quality operational demonstrations. This entire process can be summarized as follows: humans contribute intuition and diversity in operations, while agents handle optimization and stability, reducing reliance on operators, enabling execution of more complex tasks, and gathering higher-quality data.
@@ VII-B3 Sim-to-Real Paradigms
Real2Sim2real [373] enhanced imitation learning in real-world scenarios by leveraging reinforcement learning (RL) trained in a “digital twin” simulation environment. This method involves strengthening strategies through extensive RL within the simulation, followed by transferring these strategies to the real world to address data scarcity and achieve effective robotic operational imitation learning. Initially, NeRF and VR are used for scene scanning and reconstruction, and the constructed scene assets are imported into the simulator to achieve real-to-simulation fidelity. Subsequently, RL in the simulation fine-tunes the initial strategies derived from sparse expert demonstrations collected in the real world. Finally, the refined strategies are transferred to real-world settings.
@@
TRANSIC [374] narrowed the sim-to-real gap by enabling real-time human intervention to correct robot behaviors in real-world scenes. It enhances sim-to-real transfer performance through several steps: first, robots are trained using RL to establish foundational strategies within a simulation environment. Then, these strategies are implemented on real robots, with humans intervening and correcting behaviors in real-time via remote control when errors occur. The data collected from these interventions are used to train a residual policy. Integrating both foundational and residual policies ensures smoother trajectories in real-world applications following sim-to-real transfer. This approach significantly reduces the need for real-world data collection, thereby mitigating the burden while achieving effective sim-to-real transfer.
@@
Domain Randomization [375, 376, 377] enhanced the generalization of models trained in simulated environments to real-world scenarios by introducing parameter randomization during simulation. While both simulated and real environments involve perception through camera-acquired visual images, differences such as object friction and gloss make accurate simulation challenging. Randomizing parameters during simulation training covers a wide range of conditions, potentially encompassing variations that might occur in real-world settings. This approach boosts the robustness of trained models, enabling deployment from simulation to real environments.
@@
System Identification [378, 379] constructed an accurate mathematical model of physical scenes in real-world environments, encompassing parameters such as dynamics and visual rendering. It aims to make simulation environments closely resemble real-world settings, facilitating smooth transitions of models trained in simulation to real environments.
@@
Lang4sim2real [380] used natural language as a bridge to address the sim-to-real gap by using textual descriptions of images as a cross-domain unified signal. This approach aids in learning domain-invariant image representations, thereby improving generalization performance across simulation and real environments. Initially, an encoder is pretrained on image data annotated with cross-domain language descriptions. Subsequently, using the domain-invariant representations, a multi-domain, multi-task language-conditioned behavioral cloning policy is trained. This method compensates for the scarcity of real-world data by additional information from abundant simulated data, thereby enhancing sim-to-real transfer.